Calculate ΔΕ when 790 constant pressure. grams of CO2 are formed in this combustion reaction at 350 K and C10H8 (s) + 1202 (g)-10CO2 (g) + 4H20 (1) ΔΗ-5157 ki Check the states of matter in the balanced reaction. Use 44.01 g/mol as the molar mass of CO 2 2 Attempts Submit Use 5 sig figs kJ
Answer #1

In this combustion reaction, one mole C10H8 react with 12 mole O2 to form 10 mole CO2 and 4 mole H2O produced and heat released = 5157 KJ

we have

790 g CO2 are formed.

thus

moles of CO2 = mass / molar mass = 790 g / 44.01 g / mole = 17.95 mole.

thus

delta H = 5157 * 17.95 / 10 = 9256.815 KJ = 9256815 J

(delta n)g = moles of gaseous product - moles of gaseous reactant = (10 - 12) = -2

we know,

delta H = delta E + (delta n)g * RT

or

9256815 = delta E + (-2 * 8.314 * 350)

or

delta E = 9262634.8 J = 9262.6 KJ

or

delta E = 9262.6 KJ